OpenStudy (anonymous):

In a fraction with a denominator of 15, which value could be the numerator be to produce a repeating decimal A.15 B.9 C.11 D.12

OpenStudy (igreen):

11/15 = 0.733333333333333333

OpenStudy (igreen):

15/15 = 1 9/15 = 0.6 12/15 = 0.8

OpenStudy (igreen):

Therefore, your answer is C.

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @Kailamarieee

OpenStudy (triciaal):

which number when expressed as a fraction will leave a remainder? which of the numbers is a prime number?